OpenStudy (anonymous):

Suppose that 4^a=5, 5^b=6, 6^c = 7, and 7^d = 8 What is a*b*c*d?

OpenStudy (wolf1728):

4^a = 5 a * log 4 = log 5 a = log 5 / log 4 a = 0.6989700043 / 0.6020599913 a = 1.1609640474

OpenStudy (wolf1728):

5^b = 6 b * log (5) = log (6) b = log(6) / log(5) b = 1.1132827526

OpenStudy (wolf1728):

c = 1.0860331325

OpenStudy (wolf1728):

d = 1.0686215613

OpenStudy (anonymous):

\[\log_4(5)\times \log_5(6)\times\log_6(7)\times \log_7(8)\]

OpenStudy (anonymous):

if you need to find a numeric answer, you can calculate each of these using the change of base formula

OpenStudy (whpalmer4):

No need for numeric values if you observe the pattern

OpenStudy (wolf1728):

a * b * c * d = 1.1609640474 * 1.1132827526 * 1.0860331325 * 1.0686215613 = 1.5 Wow I didn't expect that

OpenStudy (anonymous):

oops i think there is more to this problem than that since the actual answer is \(\frac{3}{2}\)

OpenStudy (anonymous):

@whpalmer4 what is the gimmick here?

OpenStudy (wolf1728):

Surprising that all those calculations reduce to 1.5

OpenStudy (whpalmer4):

well, you have a = log[5]/log[4] b = log[6]/log[5] c = log[7]/log[6] d = log[8]/log[7] multiply them all together and all the intermediate ones cancel out leaving abcd = log[8]/log[4] = 1.5

OpenStudy (anonymous):

!!!

OpenStudy (whpalmer4):

yeah, whodathunkit!

OpenStudy (wolf1728):

yes I didn't bother to look for a short cut, but who got the answer ? :-)

OpenStudy (anonymous):

you, apparently

OpenStudy (wolf1728):

I guess so - and it would have been MUCH easier to reduce all those calculations

OpenStudy (whpalmer4):

think first, compute later :-)

OpenStudy (whpalmer4):

I saw the two log 5/log 4 and log 6/log 5 things you posted and the lightbulb went on...

OpenStudy (wolf1728):

Yes it's like asking what is the sum of 1 + 2 + 3 all the way to 1,000 Yes you could add 1 + 2 + 3 + 4 ... but think about it and n*(n+1) / 2 is a lot easier

OpenStudy (anonymous):

If there were a way to give two medals... :P

OpenStudy (whpalmer4):

it's the kind of problem which is just hard enough so that the brute force approach appears tractable, but takes long enough that the lazy guy might spot the pattern and look smart :-)

OpenStudy (wolf1728):

or as you said, it is the kind of problem where it seems that taking the long way is the right way